# Experience of Training a 1.7B-Parameter LLaMa Model From Scratch

### v1

@@ Abstract
Pretraining large language models is a complex endeavor influenced by multiple factors, including model architecture, data quality, training continuity, and hardware constraints. In this paper, we share insights gained from the experience of training DMaS-LLaMa-Lite, a fully open source, 1.7-billion-parameter, LLaMa-based model, on approximately 20 billion tokens of carefully curated data. We chronicle the full training trajectory, documenting how evolving validation loss levels and downstream benchmarks reflect transitions from incoherent text to fluent, contextually grounded output. Beyond standard quantitative metrics, we highlight practical considerations such as the importance of restoring optimizer states when resuming from checkpoints, and the impact of hardware changes on training stability and throughput. While qualitative evaluation provides an intuitive understanding of model improvements, our analysis extends to various performance benchmarks, demonstrating how high-quality data and thoughtful scaling enable competitive results with significantly fewer training tokens. By detailing these experiences and offering training logs, checkpoints, and sample outputs, we aim to guide future researchers and practitioners in refining their pretraining strategies. The training script is available on Github at https://github.com/McGill-DMaS/DMaS-LLaMa-Lite-Training-Code. The model checkpoints are available on Huggingface at https://huggingface.co/collections/McGill-DMaS/dmas-llama-lite-6761d97ba903f82341954ceb.
@@ I Introduction
In this paper, we present our experience pretraining DMaS-LLaMa-Lite, a 1.7-billion-parameter LLaMa-based model, on 20 billion tokens of carefully curated data. Unlike many existing studies that focus solely on the final outputs of large-scale training, we emphasize the training process itself and the insights gleaned from it. Specifically, we document the following key aspects of our work:
@@
The remainder of this paper is organized as follows: Section  II describes the model architecture, training setup, and data preprocessing. Section  III presents our findings from the pretraining experience including quantitative and qualitative evaluations and a discussion of broader implications, while Section  IV concludes with key takeaways and directions for future work.
@@ III-C Comparison and Discussion
IV Conclusion
@@
Our findings, along with the provided training logs, checkpoints, and sample outputs, aim to guide future researchers and practitioners in refining their pretraining strategies. By sharing these insights, we hope to encourage a more process-focused perspective in the development of large language models, where transparency and detailed documentation help inform design decisions, improve reproducibility, and ultimately foster more effective and efficient pretraining methodologies.

### v2

@@ Abstract
Pretraining large language models is a complex endeavor influenced by multiple factors, including model architecture, data quality, training continuity, and hardware constraints. In this paper, we share insights gained from the experience of training DMaS-LLaMa-Lite, a fully open source, 1.7-billion-parameter, LLaMa-based model, on approximately 20 billion tokens of carefully curated data. We chronicle the full training trajectory, documenting how evolving validation loss levels and downstream benchmarks reflect transitions from incoherent text to fluent, contextually grounded output. Beyond pretraining, we extend our analysis to include a post-training phase focused on instruction tuning, where the model was refined to produce more contextually appropriate, user-aligned responses. We highlight practical considerations such as the importance of restoring optimizer states when resuming from checkpoints, and the impact of hardware changes on training stability and throughput. While qualitative evaluation provides an intuitive understanding of model improvements, our analysis extends to various performance benchmarks, demonstrating how high-quality data and thoughtful scaling enable competitive results with significantly fewer training tokens. By detailing these experiences and offering training logs, checkpoints, and sample outputs, we aim to guide future researchers and practitioners in refining their pretraining strategies. The training script is available on Github at https://github.com/McGill-DMaS/DMaS-LLaMa-Lite-Training-Code. The model checkpoints are available on Huggingface at https://huggingface.co/collections/McGill-DMaS/dmas-llama-lite-6761d97ba903f82341954ceb.
@@ I Introduction
In this paper, we present our experience pretraining and post-training DMaS-LLaMa-Lite, a 1.7-billion-parameter LLaMa-based model, on 20 billion tokens of carefully curated data. Unlike many existing studies that focus solely on the final outputs of large-scale training, we emphasize the training process itself and the insights gleaned from it. Specifically, we document the following key aspects of our work:
@@
  5. 5.
    Post-Training Instruction Tuning We refine the pretrained model using parameter-efficient fintuning, leveraging high-quality instruction datasets. We documented the model’s behavioral change from the process and the model’s sensitivity to the prompt.
@@
The remainder of this paper is organized as follows: Section  II describes the model architecture, training setup, and data preprocessing. Section  III presents our findings from the pretraining experience, including quantitative and qualitative evaluations, and a discussion of broader implications. Section  IV details the post-training phase, including instruction tuning and qualitative analysis. Finally, Section  V concludes with key takeaways and directions for future work.
@@ III-C Comparison and Discussion
TABLE V: Comparison of Responses Before and After Post-Training Instruction Tuning
Prompt Before Post-Training After Post-Training
Who invented the first practical light bulb? Thomas Edison. The first practical light bulb was invented by Thomas Edison in 1879. He was a prolific inventor and held over 1,000 patents in his name…
What are the Pyramids of Giza known for? They are known for being the largest structures ever built by man. The Pyramids of Giza are known for their architectural and engineering brilliance. They were built by the ancient Egyptians as burial monuments for their pharaohs. The pyramids are considered to be the most famous and largest of the three pyramids in the Giza…
@@
IV Posttraining
Following the initial pretraining phase, we conducted a post-training step to refine the model’s capacity for following user instructions and producing contextually appropriate, helpful responses. This stage leverages techniques akin to instruction tuning, where a base model—already competent in general text generation—is further specialized to respond more effectively to prompts in a conversational format.
IV-A Instruction-Following Data and Template
For the post-training stage, we used the yahma/alpaca-cleaned  [^3] https://huggingface.co/datasets/yahma/alpaca-cleaned dataset, a widely utilized instruction-tuning corpus derived from the Stanford Alpaca dataset [22]. To present this data to the model in a consistent, conversational manner, we applied a Vicuna 1.1-style chat template [23], which structures the interaction as a user-assistant conversation:
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: input
ASSISTANT: output<|endoftext|>
Here, input is replaced with the user’s query or instruction, and output with the model’s expected answer. The ¡—endoftext—¿ token serves as a signal for the model to stop generating text at the appropriate point.
This template encourages the model to produce answers in a polite, contextually aligned, and helpful manner, while also teaching it to respect conversational boundaries. Since the training data already includes well-structured instructions and expert-crafted responses, the model can internalize the patterns and nuances of effective assistant behavior.
IV-B Model Initialization and Parameter-Efficient Fine-Tuning
We began the post-training step from the 43,500-step checkpoint of DMaS-LLaMa-Lite. Rather than fine-tuning all parameters, we employed a parameter-efficient technique based on Low-Rank Adaptation (LoRA) [24]. We injected LoRA adapters into attention-related parameters (e.g., q_proj, k_proj, v_proj, o_proj, and feed-forward projections) using a rank of $r=128$, a LoRA $\alpha=16$, and no dropout. This approach reduces the memory footprint and computational overhead while enabling efficient adaptation of the model to the instruction-following domain.
The model’s training precision and device configuration were automatically adjusted based on hardware capabilities. In this case, half-precision (FP16) or bfloat16 was used where available. Batch sizes, gradient accumulation steps, and other hyperparameters were chosen to balance computational feasibility with adequate training signal. Specifically, we set a per-device batch size of 32 and a gradient accumulation of 16 to maintain stable gradient updates, and we employed the AdamW 8-bit optimizer for memory-efficient training.
IV-C Training Dynamics
The post-training step encompassed approximately 400 steps of fine-tuning, with a linear learning rate schedule peaking at $5\times 10^{-5}$. We applied a brief warmup period of five steps to mitigate instability at the onset. During this process, the training loss steadily decreased from approximately 2.25 at the start of instruction tuning to about 1.20 by the end of the 400 steps, reflecting substantial gains in the model’s ability to generate coherent, contextually appropriate answers aligned with user instructions.
It is worth noting that the fine-tuning process does not rely on massive training resources or prolonged runs; instead, the combination of LoRA adapters and a high-quality instruction dataset enables rapid improvements. These modifications help the model internalize the format, style, and content typical of user-posed questions and the expected assistant-style responses, ultimately guiding the model toward more useful and context-sensitive outputs.
IV-D Qualitative Improvements in Responses
Table V compares the conversational behaviors of the models before and after post-training using two example prompts. Interestingly, the model demonstrates an ability to answer questions even prior to post-training. This phenomenon can be attributed to two factors. First, the Vicuna 1.1 prompting format aligns closely with natural conversational scenarios and clearly indicates that the assistant is expected to provide answers. Consequently, the model can generate responses based on its pretraining alone, without requiring additional fine-tuning. Second, it is possible that the pretraining dataset includes instruction-tuning samples formatted similarly to Vicuna 1.1 prompts, as this format predates the pretraining corpus’s knowledge cutoff.
Another noteworthy observation is that instruction tuning during post-training significantly enhances the detail and contextual depth of the model’s responses. Compared to the pretraining-only model, the post-trained version consistently generates more elaborate and contextually rich answers, reflecting improved understanding and engagement with the prompts.
IV-E Sensitivity to Prompt Variations
An intriguing phenomenon we observed is the model’s sensitivity to a small change in the prompt formatting. For example, simply appending a space after “ASSISTANT:” can alter the generated response. Without the space, the model might produce a normal answer. With the space, it often continues with a more elaborate completion:
ASSISTANT: _____________________________
The first practical light bulb was invented by Thomas Edison in 1879...
One plausible explanation involves the underlying tokenizer and how it handles spaces. In GPT-2-style tokenizers, spaces are generally prefixed to words rather than suffixed. This leads to subtle differences in the tokenization of seemingly similar inputs. For example, for an answer in the training set: ”ASSISTANT: Berlin was founded in the 13th century.”, it is tokenized as: [”ASS”, ”IST”, ”ANT”, ”:”, ”ĠBerlin”, ”Ġwas”, ”Ġfounded”, ”Ġin”, ”Ġthe”, ”Ġ13”, ”th”, ”Ġcentury”, ”.”]. Here, “Ġ” represents a space token. If the prompt ends with ”ASSISTANT: ”, however, the tokenization changes to: [”ASS”, ”IST”, ”ANT”, ”:”, ”Ġ”]. The inclusion of the extra ”Ġ” token alters the model’s understanding of the input and significantly shifts the probability distribution for the next token. This is due to the statistical co-occurrence patterns learned during pretraining, which differ depending on the presence of the additional space token.
This observation underscores the model’s inherent sensitivity to input formatting and highlights the critical role of precise prompt engineering, even down to seemingly trivial details like whitespace handling.
V Conclusion
@@
Post-training instruction tuning further refined the model’s capabilities, enabling it to respond more elaborately to a wide variety of prompts. By using parameter-efficient fine-tuning with high-quality instruction datasets, we observed significant behavioral shifts in the model. And we observed the sensitivity of the model to prompting details.
@@
Our findings, along with the provided training logs, checkpoints, and sample outputs, aim to guide future researchers and practitioners in refining both pretraining and post-training strategies. By documenting the full trajectory from pretraining dynamics to post-training refinement, we emphasize a process-focused perspective where transparency and detailed insights foster reproducibility and more effective model development. This approach lays a foundation for future studies to explore additional best practices, methodological refinements, and novel techniques in both pretraining and instruction-tuning phases of large language models.